Joni Received $25,000 as part of a settlement in a class action lawsuit. She invested some money at 10% and the rest at a 9% simple interest rate. If her total annual income from these two investments was $2,430, how much did she invest at each rate?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

Part I 10.00% per annum ------------- Amount invested =x
Part II 9.00% per annum ------------ Amount invested = y
25000
Interest----- 2340.00

Part I 10.00% per annum ---x
Part II 9.00% per annum ---y
Total investment
x + 1 y= 25000 -------------1
Interest on both investments
10.00% x + 9.00% y= 2340
Multiply by 100
10 x + 9 y= 234000.00 --------2
Multiply (1) by -10
we get
-10 x -10 y= -250000.00
Add this to (2)
0 x -1 y= -16000
divide by -1
y = 16000
Part I 10.00% $ 9000
Part II 9.00% $ 16000

CHECK
9000 --------- 10.00% ------- 900.00
16000 ------------- 9.00% ------- 1440.00
Total -------------------- 2340.00

m.ananth@hotmail.ca


Answer by richwmiller(17219)

x+y=25000,
0.1*x+0.09*y=2430
x=25000-y
0.1*(25000-y)+0.09*y=2430
2500-0.1y+0.09*y=2430
-0.01*y=-70
y=7000 at 9%
x=25000-y
x=18000 at 10%
check
0.1*18000+0.09*7000=2430
1800+630=2430
2430=2430
ok